Find the condition that the point (x,y) lies on the line joining (3,4) and (-3,-6)

Grade:10

1 Answers

Arun
6 years ago
Dear Palak
 
for this, we have to find the equation of line using (x,y)
hence
 
y – 4 = [(-6 – 4)/(-3 – 3) ] *(x –3)
 
y – 4 = (5/3) (x – 3)
3 y – 12 = 5x – 15
5x – 3y = 3
 
hence (x, y) should satisfy this condition
 
Regards
Arun (askIITians forum expert)